The mass of mars 0.11 times the mass of the earth and radius of 0.532 times of the earth. At what minimum velocity should a rocket be released from the surface of the mars so that it can come out of the gravitational force of the mars?


1
Expert's answer
2021-06-29T08:18:03-0400
"V_e=\sqrt{\frac{2GM}{r}}\\
V_M=\sqrt{\frac{M_mr_e}{r_mM_e}}V_E
\\
V_M=\sqrt{\frac{0.11}{0.532}}11.2=5.1\frac{km}{s}"